OpenStudy (anonymous):

factorise e'2' - 100?

OpenStudy (lalaly):

difference between two squares x^2-a^2 = (x-a)(x+a)

Parth (parthkohli):

This is a difference of squares. \(\Large \color{MidnightBlue}{\Rightarrow a^2 - b^2 = (a + b)(a - b) }\) Hint, 100 = 10^2

OpenStudy (anonymous):

so whats the answer

Parth (parthkohli):

Wow -.-

Parth (parthkohli):

We just taught it to you.

OpenStudy (lalaly):

We dont give answers, u find the answer and we will tell u if its right or wrong

Parth (parthkohli):

I think he's gone

OpenStudy (lalaly):

oh well, too bad lol

Parth (parthkohli):

hehe